OpenStudy (anonymous):

Write down the first five terms of each sequence {sn}= {n^2+1}

OpenStudy (anonymous):

replace n by 1,2,3, 4, 5 get \[s_1=1^2+1=2\] \[s_2=2^2+1=5\] \[s_3=3^2+1=10\] etc

OpenStudy (anonymous):

do you know how to do this kind? Write out the first five terms of the sequence: a1 = 3 an = 4 - a(n -1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[a_1=3\] \[a_2=4-a_{2-1}=4-a_1=4-3=1\]

OpenStudy (anonymous):

\[a_3=4=a_2=4-1=3\]

OpenStudy (anonymous):

\[a_4=4-3=1\] so it looks like you just get 4 and 1

OpenStudy (anonymous):

odd example for a recursive formula i must say.